# Barycenter curvature-dimension condition for extended metric measure spaces

### v1

@@ 3 Wasserstein barycenter
Example 3.5 (See [EH15]).
Let $(\mathbf{Y},\mathsf{d}_{\mathbf{Y}},\pi)$ be an extended metric measure space consisting of the configuration space $\mathbf{Y}$ over a Riemannian manifold $M$, the intrinsic metric $\mathsf{d}_{\mathbf{Y}}$ and the Poisson measure $\pi$. Assume that $M$ has Ricci curvature bounded below by $K\in\mathbb{R}$.
Then for any probability measure $\Omega$ over $\mathcal{P}(\mathbf{Y})$ with finite variance, its barycenter $\bar{\mu}$ satisfies
${\rm Ent}_{\pi}(\bar{\mu})\leq\int_{\mathcal{P}(X)}{\rm Ent}_{\pi}(\mu)\,{\mathrm{d}}\Omega(\mu)-\frac{K}{2}\int_{\mathcal{P}(X)}W_{2}^{2}(\bar{\mu},\mu)\,{\mathrm{d}}\Omega(\mu).$
@@ 4 Barycenter Curvature-Dimension condition
We have seen the links between displacement convexity, Wasserstein barycenter (Jensen’s inequality) and curvature-dimension conditions of metric measure spaces. This motivates us to propose a new curvature-dimension condition, which covers some non-trivial extended metric measure spaces (also non-geodesic), e.g. Wiener spaces and Configuration spaces over a Riemannian manifold.
Since barycenter plays a crucial role, we call this new curvature-dimension condition, Barycenter Curvature-Dimension condition and denote it by BCD.
@@
If we take $\Omega=(1-t)\delta_{\mu_{0}}+t\delta_{\mu_{1}}$ and $(X,\mathsf{d})$ is a geodesic space, (4.1) implies the Lott–Sturm–Villani curvature-dimension condition.
@@
Similar to Lott–Sturm–Villani’s ${\rm CD}(K,N)$ condition (cf. [LV09, Stu06b]) and Bakry–Émery’s ${\rm BE}(K,N)$ (cf. [BGL14]), there is a finite dimensional version of $\rm BCD$ condition, ${\rm BCD}(K,N)$ condition; see [HLZ24] for details.
Not surprisingly, we can establish that the moduli space of compact metric measure spaces satisfying barycenter curvature-dimension condition, is closed under measured Gromov–Hausdorff convergence.
@@
As explained in Section 2, the difference between $\rm CD$ condition and $\rm RCD$ condition lies in whether they have a “Riemannian structure” or not. It is natural to ask:
Problem 1:
Do Finsler manifolds with lower Ricci curvature bound from below satisfy BCD condition? Is there a “Riemmanian Barycenter-curvature dimension condition”?
Problem 2:
What is the difference between $\rm CD$ condition and $\rm BCD$ condition?
Since spaces satisfying $\rm BCD$ condition, we have discovered, are geodesic or “almost geodesic”, it is interesting to find a $\rm BCD$ space which is not geodesic.
Problem 3:
Is there a non-geodesic $\rm BCD$ space?
At the end of this paper, we introduce some new geometric inequalities, as simple but interesting applications of our BCD theory. For simplicity, we will only deal with ${\rm BCD}$ metric measure spaces. We also refer the readers to a recent paper [KW22] by Kolesnikov and Werner for more applications of Wasserstein barycenter.
@@
$E:=\left\{x~{}\text{is the barycenter of}~{}\sum_{i=1}^{n}\lambda_{i}\delta_{x_{i}}:x_{i}\in E_{i},i=1,\dots,n\right\}.$

### v2

@@ 3 Wasserstein barycenter
Furthermore, if $\int_{\mathcal{P}(X)}{\rm Ent}_{\gamma}(\mu)\,{\mathrm{d}}\Omega(\mu)<+\infty$, $\Omega$ has a unique barycenter.
Corollary 3.6 (See [EH15]).
Let $(\mathbf{Y},\mathsf{d}_{\mathbf{Y}},\pi)$ be an extended metric measure space where $\mathbf{Y}$ is the configuration space over a Riemannian manifold $M$, $\mathsf{d}_{\mathbf{Y}}$ is the intrinsic metric and $\pi$ is the Poisson measure . Assume that $M$ has Ricci curvature bounded from below by $K\in\mathbb{R}$.
Then for any barycenter $\bar{\mu}$ of a probability measure $\Omega$ over $\mathcal{P}(\mathbf{Y})$ with finite variance, its holds
${\rm Ent}_{\pi}(\bar{\mu})\leq\int_{\mathcal{P}(\mathbf{Y})}{\rm Ent}_{\pi}(\mu)\,{\mathrm{d}}\Omega(\mu)-\frac{K}{2}\int_{\mathcal{P}(\mathbf{Y})}W_{2}^{2}(\bar{\mu},\mu)\,{\mathrm{d}}\Omega(\mu).$
Furthermore, if $\int_{\mathcal{P}(\mathbf{Y})}{\rm Ent}_{\pi}(\mu)\,{\mathrm{d}}\Omega(\mu)<+\infty$, $\Omega$ has a unique barycenter.
@@ 4 Barycenter Curvature-Dimension condition
In the previous sections, we have seen the connection between displacement convexity, Wasserstein barycenter, Jensen’s inequality and curvature-dimension condition of metric measure spaces. This motivates us to propose a new curvature-dimension condition, which covers some non-trivial extended metric measure spaces, e.g. Wiener spaces and Configuration spaces over a Riemannian manifold.
Due to the special role of barycenter, we call this new curvature-dimension condition Barycenter Curvature-Dimension condition and denote it by BCD for short.
@@
If we take $\Omega=(1-t)\delta_{\mu_{0}}+t\delta_{\mu_{1}}$ and $(X,\mathsf{d})$ is a geodesic space, BCD condition implies the Lott–Sturm–Villani’s curvature-dimension condition.
@@
Similar to ${\rm CD}(K,N)$ condition (cf. [LV09, Stu06b]) and ${\rm RCD}(K,N)$ condition (cf. [EKS15]), there is a finite dimensional version of $\rm BCD$ condition, called ${\rm BCD}(K,N)$ condition; see [HLZ24, Section 6] for details.
Not surprisingly, we can prove that the family of compact metric measure spaces satisfying barycenter curvature-dimension condition, is closed in measured Gromov–Hausdorff topology.
@@
Next we introduce some new geometric inequalities, as simple but interesting applications of our BCD theory. For simplicity, we will only deal with metric measure spaces. We refer the readers to a recent paper [KW22] by Kolesnikov and Werner for more applications of Wasserstein barycenter. In addition, see [HLZ24] for more applications of our BCD theory to multi-marginal optimal transport.
@@
$E:=\left\{x~{}\text{is a barycenter of}~{}\sum_{i=1}^{n}\lambda_{i}\delta_{x_{i}}:x_{i}\in E_{i},i=1,\dots,n\right\}.$
@@
At the end of this paper, we list some problems we would like to study in the future.
Open Problem 4.7.
As explained in Section 2, the difference between $\rm CD$ condition and $\rm RCD$ condition lies in whether they have a “Riemannian structure” or not. It is natural to ask: Do Finsler manifolds with lower Ricci curvature bound from below satisfy BCD condition? Is there a “Riemmanian Barycenter-curvature dimension condition”?
Open Problem 4.8.
Can we find more similarities and differences between $\rm CD$ condition and $\rm BCD$ condition?
Open Problem 4.9.
In the preceding examples of BCD spaces, they are either geodesic or “almost geodesic”. It is interesting to find a $\rm BCD$ space with less geodesics. Moreover, is there a $\rm BCD$ space which has a fractal structure?